# Chronological Passage Assembling in RAG framework for Temporal Question Answering

### v1

@@ Abstract
Long-context question answering over narrative tasks is challenging because correct answers often hinge on reconstructing a coherent timeline of events while preserving contextual flow in a limited context window. Retrieval-augmented generation (RAG) indexing methods aim to address this challenge by selectively retrieving only necessary document segments. However, narrative texts possess unique characteristics that limit the effectiveness of these existing approaches. Specifically, understanding narrative texts requires more than isolated segments, as the broader context and sequential relationships between segments are crucial for comprehension. To address these limitations, we propose ChronoRAG, a novel RAG framework specialized for narrative texts. This approach focuses on two essential aspects: refining dispersed document information into coherent and structured passages, and preserving narrative flow by explicitly capturing and maintaining the temporal order among retrieved passages. We empirically demonstrate the effectiveness of ChronoRAG through experiments on the NarrativeQA dataset, showing substantial improvements in tasks requiring both factual identification and comprehension of complex sequential relationships, underscoring that reasoning over temporal order is crucial in resolving narrative QA.
@@ 1 Introduction
Figure 1 illustrates the mismatch between conventional retrieval strategies and the characteristics of narrative data. A common approach, as shown in (a) of Figure 1, is to retrieve as many sentences as possible that are likely to match the query, often based on textual similarity. To do so, documents are typically stored as isolated sentences. While such methods may successfully retrieve a sentence containing the correct answer, they often fail to provide sufficient contextual cues. For instance, this can create ambiguity, making it unclear whether "London" or "Paris" is the location relevant to the question, even if both are mentioned in the retrieved results.
@@
We empirically validate our proposed approach on the NarrativeQA Kočiskỳ et al. (2018) dataset. To rigorously test temporal reasoning, we isolate a subset of "Time Questions" that require understanding event sequences. Our experiments show that our method achieves significant improvements in both the complete dataset and the specialized Time Question set. Notably, these results are achieved using lighter graph construction and retrieval mechanisms than those found in existing summary and graph-based methods, demonstrating enhanced performance in identifying individual facts and comprehending complex relational structures.
@@
  • Experiments on the NarrativeQA dataset demonstrate the effectiveness of our framework, and emphasizing event-to-event relations drives performance gains for both factual and temporal queries, highlighting the critical role of relational understanding over entity extraction.
@@ 2 Related Work
Passage Granularity. Document indexing approaches have been explored with varying passage granularity to improve retrieval precision. DenseXRetrieval Chen et al. (2024) advocates finer granularities to enhance information precision. Conversely, MolecularFacts Gunjal and Durrett (2024) demonstrates that overly granular decompositions such as atomic facts or propositions often lose critical contextual cues, advocating instead for concise yet contextually coherent units. Our method employs atomic facts as keys for retrieval while preserving broader narrative flows as the retrieved values.
Summary-Based Document Augmentation. Summary-based indexing methods, including RAPTOR Sarthi et al. (2024), MemWalker Chen et al. (2023), and ReadAgent Lee et al. (2024) leverage iterative summarization to build hierarchical structures that improve retrieval accuracy and contextual coherence. However, such methods often suffer from high computational costs due to deep hierarchical structures and redundant overlapping information. Our approach simplifies the hierarchical concept by adopting a single-layer summary, significantly reducing computation and overlap issues while maintaining contextual effectiveness.
Knowledge Graph-Based Document Augmentation. Graph-based augmentation methods, such as GraphRAG Edge et al. (2024) and LightRAG Guo et al. (2024), typically construct knowledge graphs by extracting entities and relationships from documents. These methods are good at capturing entity-centric information but struggle to represent relationships between entities explicitly, a crucial element in narratives. Our proposed framework explicitly incorporates sequential relations among narrative elements, addressing this critical limitation of traditional knowledge graphs.
@@ 3 ChronoRAG
Figure 2: Overall Pipeline of ChronoRAG.
@@
We present ChronoRAG, a novel Retrieval-Augmented Generation (RAG) framework specialized for narrative texts where chronological context is crucial. As described in Figure 2, our framework is composed of two primary stages: an offline Graph Construction phase where the original documents are processed into a hierarchical, linked structure, and an online Passage Retrieval and Answer Generation phase where the constructed graph is used to answer queries.
3.1 Graph Construction
The goal of this offline phase is to transform a raw document into a structured, two-layer graph that captures both factual information and narrative chronology. This process involves the following steps:
Document Chunking. Due to inherent limitations in handling entire documents simultaneously, we first divide the original document into fixed-length chunks (e.g., 100 tokens each). These chunks constitute Layer 0 of our hierarchical graph, facilitating consistent and manageable retrieval. While fixed-length chunking may disrupt internal narrative coherence, maintaining a consistent token length is crucial for retrieval. Hence, we avoid semantic chunking with variable lengths.
Summarizing Chunks. Next, we group chunks (e.g., 10 chunks per group) and summarize them by instructing LLM. This summarization distills complex narrative passages, clarifying overall content and creating more manageable retrieval units.
Entity-Relation Extraction. Then we instruct the LLM to extract entities from summarized texts and generate relational descriptions between entities. We utilize only relation descriptions for indexing and retrieval, avoiding overlapping entity descriptions that might disrupt narrative flow. These relations, functioning like atomic facts, constitute Layer 1 in the hierarchical graph, enhancing retrieval precision due to their focused and coherent informational structure.
Indexing. We assign narrative-order indices to both relation description sentences derived from summaries and to original document chunks. Relation descriptions are indexed according to the position of their source chunks in the document and the order in which the descriptions were generated, with lower indices assigned to those derived from earlier chunks or generated earlier in the process. Furthermore, each relation sentence stores the index of the original chunk it was derived from as a child index, enabling quick access to neighboring relations and original chunks.
Neighborhood Assembling. We augment retrieved relational descriptions with their surrounding context to reconstruct a narrative flow. Rather than relying on isolated facts, we aim to provide contextually rich information. Specifically, for Layer 1 retrievals, we separate the role of the retrieved item into a key (the relation description itself) and a value (neighboring Layer 1 passages concatenated in index order). This approach allows us to preserve a coherent local storyline rather than referencing fragmented facts.
3.2 Retrieving Passage
@@
At inference time, a query is handled through a hierarchical retrieval process that leverages the constructed graph to assemble a rich, chronologically-aware context for the LLM.
@@
Hierarchical Retrieving. We leverage the hierarchical granularity of Layer 1 and Layer 0 for retrieval. We begin by retrieving high-precision relation descriptions from Layer 1. Then, using the associated child indices, we retrieve related Layer 0 chunks, ensuring a comprehensive and balanced context. This is crucial because Layer 0 often retains omitted details and original dialogues that are valuable for question answering.
@@
Answer Generation. We combine the original query with the context obtained through hierarchical retrieval and feed them into the language model. Each passage is separated by double line breaks and ordered by relevance, enabling accurate and coherent answer generation.
@@ 4.1 Experimental Setup
Dataset. We employ the NarrativeQA Kočiskỳ et al. (2018) as our primary dataset. NarrativeQA comprises 355 stories and scripts with a total of 10,557 question–answer pairs. From this pool, we identify and separate a subset of 1,111 Time Questions, defined as those containing temporal keywords {“When,” “While,” “During,” “After,” “Before”}. These Time Questions require retrieving and reasoning over multiple related events, making them a particularly challenging subset for temporal understanding and reasoning.
Evaluation Metric. We measure answer quality using ROUGE-L Lin (2004), which computes the Longest Common Subsequence overlap between a generated answer and its corresponding human reference. Due to the short and pronoun-heavy nature of NarrativeQA answers, ROUGE-L effectively captures agreement in key word sequences without penalizing minor rephrasings.
Baselines. We compare against five existing methods that differ in information extraction, representation, and retrieval structure:
@@
Implementation Details. All baselines share identical hyperparameter settings: top-k of 20 for retrieval, contextTokenLengthLimit of 1,500 tokens, and the same sampling strategy during generation. We perform all summarization and entity–relation extraction steps with meta-llama-3-8B-Instruct Grattafiori et al. (2024). We compute retrieval scores using embedding similarity exclusively; we don’t use BM25 Robertson et al. (2009) to prevent distortion of the original text during generation. Specifically, we employ the arctic-Snowflake-embed-l Merrick et al. (2024) for generating embeddings, and use unifiedqa-v2-t5-3b-1363200 Khashabi et al. (2022) for final answer generation. All retrieved contexts fed into the generator respect the 1,500-token length limit to ensure fair comparison across methods.
@@ 4.2 Main Results
Performance Comparison. Table 1 shows that ChronoRAG, our proposed approach outperforms on both the full NarrativeQA full dataset and the Time Question subset compared to baselines. Summarization based baselines such as RAPTOR-CT and RAPTOR-TT follow, but they still lag. The results indicate that restructuring events into a clear temporal order supplies the language model with the most coherent context for narrative reasoning and question answering. GraphRAG records the lowest score among all methods for several reasons. Its exhaustive entity–relation extraction adds thousands of trivial nodes, inflating the graph and burying key plot elements under noise. And it omits the covariate filtering stage, so graph expansion begins from noisy entity-relations and quickly drifts into irrelevant subgraphs. These compounded issues dilute precision so severely that GraphRAG scores lowest.
@@
NaiveRAG 0.255 0.227
Propositionizer 0.262 0.238
RAPTOR_CT 0.297 0.261
RAPTOR_TT 0.295 0.259
LightRAG 0.240 0.214
GraphRAG 0.200 0.185
ChronoRAG 0.308 0.268
Table 1: QA Performance on NarrativeQA (ROUGE-L).
Top performance is bolded, Second best is underlined.
Ablation Study. In this section, we conduct ablation studies to investigate the effectiveness of different components and settings of ChronoRAG. As shown in the Table 2, without summarizing the original text and extracting entity relations shows a significant performance degradation, showing the importance of chunk summarization. The effects of summarization are twofold: it leaves only important information, making retrieval easier, and when assembling, it clarifies the flow. The results without passage assembling are obtained by individually searching for entity relations extracted from the summary, while the results without chunk summarization are obtained by searching for entity relations directly extracted from the 10 chunks. Despite not connecting nearby passages in both settings, a significant performance difference is observed in the TimeQuestion.
Ablation Whole Data Time Question
@@
Table 2: Ablation Study on NarrativeQA (ROUGE-L)
Trade-off between Linking Window and the Number of Retrieved Passage. While extending the connection beyond adjacent text segments can enhance the local contextual coherence of retrieved passages, it also increases the length of each segment, thereby reducing the total number of passages that can be retrieved within a fixed token limit. Our experiments revealed that this trade-off has a detrimental effect on retrieval quality. when the number of adjacent sentences included in each passage was increased to two, overall and temporal questions performance declined to 0.300 and 0.258, respectively.
Computation Costs. Our pipeline is computationally efficient, requiring just two LLM calls per 1,000 tokens for graph construction. Although this cost increases linearly with document length, it remains lower than competing methods like recursive summarization. Furthermore, only one LLM call is required for answer generation during search, with our method still attaining the highest performance despite its efficiency.
Case Study.
Figure 3: Retrieved context per method.

### v2

@@ Abstract
Long-context question answering over narrative tasks is challenging because correct answers often hinge on reconstructing a coherent timeline of events while preserving contextual flow in a limited context window. Retrieval-augmented generation (RAG) methods aim to address this challenge by selectively retrieving only necessary document segments. However, narrative texts possess unique characteristics that limit the effectiveness of these existing approaches. Specifically, understanding narrative texts requires more than isolated segments, as the broader context and sequential relationships between segments are crucial for comprehension. To address these limitations, we propose ChronoRAG, a novel RAG framework specialized for narrative texts. This approach focuses on two essential aspects: refining dispersed document information into coherent and structured passages and preserving narrative flow by explicitly capturing and maintaining the temporal order among retrieved passages. We empirically demonstrate the effectiveness of ChronoRAG through experiments on the NarrativeQA and GutenQA dataset, showing substantial improvements in tasks requiring both factual identification and comprehension of complex sequential relationships, underscoring that reasoning over temporal order is crucial in resolving narrative QA.  [^1] The source code will be released upon paper acceptance.
@@ 1 Introduction
Figure 1 illustrates the mismatch between conventional retrieval strategies and the characteristics of narrative data. As shown in (a) of Figure 1, a common approach is to retrieve as many sentences as possible that are likely to match the query based on textual similarity. To do so, documents are typically stored as isolated sentences. While such methods may successfully retrieve a sentence containing the correct answer, they often fail to provide sufficient contextual cues. This can create ambiguity, making it unclear whether "London" or "Paris" is the location relevant to the question, even if both are mentioned in the retrieved results.
@@
We empirically validate our proposed approach on the NarrativeQA Kočiskỳ et al. (2018) and GutenQA Duarte et al. (2024). To rigorously test temporal reasoning, we isolate a subset of "Time Questions" that require understanding event sequences. Our experiments show that our method achieves significant improvements in both the complete dataset and the specialized Time Question set. Notably, these results are achieved using lighter graph construction and retrieval mechanisms than those found in existing summary and graph-based methods, demonstrating enhanced performance in identifying individual facts and comprehending complex relational structures.
Our contributions can be summarized as follows:
@@
  • Experiments demonstrate the effectiveness of our framework, and emphasizing event-to-event relations drives performance gains for both factual and temporal queries, highlighting the critical role of relational understanding over entity extraction.
@@ 2 Related Work
Passage Granularity Document indexing approaches have been explored with varying passage granularity to improve retrieval precision. DenseXRetrieval Chen et al. (2024) advocates finer granularities to enhance information precision. Conversely, MolecularFacts Gunjal and Durrett (2024) demonstrates that overly granular decompositions such as atomic facts or propositions often lose critical contextual cues, advocating instead for concise yet contextually coherent units. Our method strikes a balance by using atomic facts as keys for indexing, while preserving broader narrative flows as values during retrieval, thereby combining precision with coherence.
Summary-Based Document Augmentation
Summary-based indexing methods construct hierarchical structures by iteratively compressing document segments into progressively shorter summaries. RAPTOR Sarthi et al. (2024), MemWalker Chen et al. (2023), and ReadAgent Lee et al. (2024) demonstrate how such designs enhance retrieval accuracy and contextual coherence by filtering out irrelevant passages and aggregating query-focused summaries. However, the deep hierarchies adopted in these approaches often lead to redundant overlaps and high computational costs. Our approach simplifies the hierarchical concept by adopting a single-layer summary, significantly reducing computation and overlap issues while maintaining contextual effectiveness.
Figure 2: The offline Graph Construction pipeline of ChronoRAG. This process transforms an unstructured narrative document into a structured, two-layer graph that explicitly encodes chronological relationships.
Knowledge Graph-Based Document Augmentation
Graph-based augmentation represents another line of work that emphasizes structured relational knowledge. GraphRAG Edge et al. (2024) formalizes the paradigm through components such as query processors, retrievers, organizers, and generators, and leverages graph traversal and community detection to retrieve information beyond lexical similarity. LightRAG Guo et al. (2024) reduces preprocessing and latency overhead by encoding relational signals into dense indices and employing coarse-to-fine retrieval. Extensions such as EventRAG Yang et al. (2025b) construct event knowledge graphs that capture temporal and causal dependencies, while Entity–Event RAG Zhang et al. (2025) prevents collapsing distinct entity mentions by maintaining separate but linked entity and event subgraphs. Iterative reasoning approaches, such as KG-IRAG Yang et al. (2025a), further refine this idea by incrementally retrieving over temporal and logical dependencies. Despite these advances, most methods emphasize static entity-centric relations; our framework differs by explicitly modeling sequential narrative relations, thereby addressing a critical gap in capturing dynamic contextual flows.
@@ 3 ChronoRAG
Figure 3: The online Passage Retrieval process of ChronoRAG for a sample query. This demonstrates how the pre-constructed graph is used at inference time to assemble a chronologically coherent context for the LLM.
@@
We present ChronoRAG, a novel RAG framework specialized for narrative texts where chronological context is crucial. Most RAG systems treat documents as a collection of independent facts, which fragments the timeline and severs the contextual links essential for understanding sequential events. To address this, we design our framework to reconstruct narrative flow by explicitly modeling and preserving the temporal order of events. As described in Figure 2, our framework is composed of two primary stages: an offline Graph Construction phase where the original documents are processed into a hierarchical, linked structure, and an online Passage Retrieval and Answer Generation phase where the constructed graph is used to answer queries.
3.1 Offline Graph Construction
This offline phase transforms a raw document into a structured, two-layer graph that captures both factual information and narrative chronology. As shown in Figure 2, this graph is composed of two levels: a foundational Layer 0 containing the original document text divided into sequential, fixed-length chunks, which preserves narrative detail, and an abstract Layer 1 built from concise, structured relation descriptions that represent the key events and relationships. The graph construction process involves four steps.
3.1.1 Document Chunking
Due to inherent limitations in processing an entire document simultaneously, we first divide the original document ($D$) into fixed-length chunks ($d_{i}$), each consisting of up to $k$ tokens. This approach ensures that all retrieved document segments fit within a predefined context length, thereby maintaining both the quality and manageability of retrieval results. The document is initially segmented into individual sentences, which are then sequentially appended to each chunk. When the cumulative length exceeds $k$ tokens, the next sentence is assigned to a new chunk, ensuring that sentences are not split across chunks. In rare cases where a single sentence itself exceeds $k$ tokens, the sentence is split to guarantee that every chunk remains within the token limit. These segmented document chunks serve as the fundamental retrieval units and constitute Layer 0 of the graph constructed in subsequent stages. The document $D$ is thus represented as a set of chunks, where each chunk $d_{i}$ has a token count less than $k$, as follows:
$D=\{d_{1},d_{2},d_{3},\ldots,d_{i}\},\quad|d_{i}|<k$
3.1.2 Chunk Summarization
Next, we sequentially cluster the chunks in the original document order, grouping every $l$ chunks. We concatenate and summarize each cluster’s texts using an LLM. We utilize sequential clustering because it preserves the document’s original order while maintaining a controllable and consistent input length for the LLM. This summarization step facilitates higher-level representation learning by focusing on the overall flow of the document rather than retaining excessive local detail. For example, the left panel of Figure 2 shows how a raw Document Chunk about Madeline Hall is condensed into a more concise Chunk Summary. For each cluster of $l$ chunks, we generate a summary $S_{i}$ by an LLM using a summarization prompt $P_{\mathrm{summarize}}$ as formulated below: (We provide a full prompt in Appendix C.)
$s_{i}=\mathrm{LLM}(P_{\mathrm{summarize}},\{d_{i1},d_{i2},d_{i3},\ldots,d_{il}\})$
3.1.3 Entity-Relation Extraction
We then transform the generated summaries via LLM into relational descriptions among entities. We adapt the prompt of GraphRAG Edge et al. (2024) into a one-shot instruction for entity–relation extraction. (Full prompts are in Table 6 of Appendix.) From LLM’s outputs, which consist of both entity descriptions($E_{i}$) and relation descriptions($R_{i}$), we only use the relation descriptions to form the Layer 1 nodes of the graph. This extraction step decomposes the summarized text into retrieval-friendly fragments, as described in the left panel of Figure 2, where the summary is broken down into atomic facts like "Madeline Hall is owned by Miss Delmar".
Here, we exclude entity descriptions because identical entities may appear redundantly across multiple chunks. Finally, we formalize this extraction step, where an LLM processes each summary $S_{i}$ to produce a set of entities $E_{i}$ and relations $R_{i}$, as shown below: (Full prompts are in Appendix C.)
$\{E_{i},R_{i}\}=\mathrm{LLM}(P_{\mathrm{extraction}},S_{i})$
3.1.4 Hierarchical and Temporal Indexing
This step involves assigning indices to the relation description sentences ($R_{i}$) derived from the summary and the document chunks ($d_{i}$) from the original text. Document chunks are indexed sequentially according to their original order in the source text, while the relation description sentences are indexed either based on the earlier chunks from which they are derived and in the order in which they were generated during information extraction.
Each document chunk corresponds to a Layer-0 node, and each relation sentence forms a Layer-1 node. For quick access, each Layer-1 node connects to its corresponding Layer-0 nodes—those within the cluster from which it was derived—by establishing directed edges. Additionally, adjacent Layer-1 nodes (according to their index) are also linked via edges. This indexing and edge-construction process results in the formation of a unified graph structure.
3.2 Online Passage Retrieval
@@
At inference time, we handle a query through a hierarchical retrieval process that leverages the constructed graph to assemble a rich, chronologically-aware context for the LLM.
3.2.1 Hierarchical Retrieving
@@
We leverage the hierarchical granularity of Layer 1 and Layer 0 for retrieval. We begin by retrieving high-precision relation descriptions from Layer 1 based on semantic similarity to the query. Then, using the links established during indexing, we retrieve the related Layer 0 chunks to provide a comprehensive and balanced context. As illustrated in Figure 3, this process first identifies a key event in Layer 1 and later retrieves the detailed source text from Layer 0. This step is crucial because Layer 0 often retains omitted details and original dialogues that are valuable for question answering.
3.2.2 Neighborhood Assembling
We then augment retrieved relational descriptions with their surrounding context to reconstruct a narrative flow. Rather than relying on isolated facts, we aim to provide contextually rich information. As in the example of Figure 3, after ChronoRAG retrieves a key event, such as "Gandalf orders Aragorn to lead the Fellowship of the Ring (idx 39)", the system automatically appends its chronological neighbors, including "Gandalf confronts the Balrog... (idx 38)" and "Gandalf falls off a cliff... (idx 41)". This creates a coherent, temporally ordered passage that preserves the local storyline, providing crucial context that isolated facts would lack.
3.2.3 Answer Generation
@@
Finally, we combine the original query with the context obtained through hierarchical retrieval and neighborhood assembling and feed them into the language model. We separate each passage by double line breaks and sort by relevance, enabling accurate and coherent answer generation.
@@ 4.1 Experimental Setup
Dataset
We employ the NarrativeQA Kočiskỳ et al. (2018) and GutenQA Duarte et al. (2024) datasets to measure ChronoRAG’s performance. NarrativeQA comprises 10,557 question-answer pairs from 391 stories, while GutenQA consists of 3,000 pairs from 1,000 stories. To specifically evaluate temporal reasoning, we construct a "Time Questions" subset by selecting all samples containing at least one of a set of predefined temporal keywords: {‘When,’ ‘While,’ ‘During,’ ‘After,’ ‘Before’}. This process yields 1,111 questions from NarrativeQA and 662 from GutenQA, respectively. These questions require retrieving and reasoning over multiple related events, making them a demanding benchmark for temporal understanding.
NarrativeQA
Metric ROUGE CosineSim LLM Eval (ACC)
Subset Whole Data Time Question Whole Data Time Question Whole Data Time Question
NaiveRAG 0.255 0.227 0.841 0.844 0.183 0.144
Propositionizer 0.262 0.238 0.846 0.852 0.189 0.141
RAPTOR_CT 0.298 0.262 0.854 0.858 0.241 0.178
RAPTOR_TT 0.289 0.253 0.851 0.854 0.231 0.170
LightRAG 0.240 0.214 0.841 0.845 0.182 0.123
GraphRAG 0.195 0.185 0.823 0.830 0.139 0.106
ChronoRAG (Ours) 0.308 0.268 0.853 0.854 0.257 0.195
GutenQA
Metric ROUGE CosineSim LLM Eval (ACC)
Subset Whole Data Time Question Whole Data Time Question Whole Data Time Question
NaiveRAG 0.166 0.172 0.769 0.778 0.251 0.278
Propositionizer 0.151 0.142 0.779 0.785 0.167 0.140
RAPTOR_CT 0.159 0.164 0.775 0.784 0.244 0.269
RAPTOR_TT 0.104 0.102 0.762 0.769 0.134 0.119
LightRAG 0.083 0.083 0.740 0.750 0.075 0.077
GraphRAG 0.122 0.115 0.760 0.767 0.134 0.119
ChronoRAG (Ours) 0.159 0.170 0.776 0.785 0.248 0.275
Table 1: QA Performance on NarrativeQA and GutenQA across ROUGE, CosineSim, and LLM Eval metrics. Top performance is bolded, Second best is underlined.
Evaluation Metric
We measure answer quality using ROUGE-L Lin (2004), which computes the Longest Common Subsequence (LCS) overlap between a generated answer and its corresponding human reference. Due to the short and pronoun-heavy nature of NarrativeQA answers, ROUGE-L effectively captures agreement in key word sequences without penalizing minor rephrasings.
In addition to ROUGE-L, we employ cosine similarity and LLM-based evaluation to better assess semantic fidelity. ROUGE may fail to capture semantically similar answers that differ lexically, so we incorporate complementary metrics to address this limitation. We compute cosine similarity using the Snowflake model Merrick et al. (2024), which is also used during the retrieval process. It measures the embedding-based similarity between the generated answer and the reference answer.
For LLM-based evaluation, we use GPT-4.1 Mini Achiam et al. (2023). Given the question, summary, gold passage, and ground-truth answer, the GPT model performs binary classification—[Correct] or [Wrong]—to determine whether the generated answer can be considered valid. See Appendix C for the detailed prompt.
Baselines
We compare against five existing methods that differ in information extraction, representation, and retrieval structure:
@@
Implementation Details.
All baselines share identical hyperparameter settings: top-k of 20 for retrieval, contextTokenLengthLimit of 1,500 tokens, and the same greedy decoding strategy during generation. We perform all summarization and entity–relation extraction steps with meta-llama-3-8B-Instruct Grattafiori et al. (2024). We compute retrieval scores using embedding similarity exclusively; we don’t use BM25 Robertson et al. (2009) to prevent distortion of the original text during generation. Specifically, we employ the arctic-Snowflake-embed-l Merrick et al. (2024) for generating embeddings, and use unifiedqa-v2-t5-3b-1363200 Khashabi et al. (2022) for final answer generation. All retrieved contexts fed into the generator respect the 1,500-token length limit to ensure fair comparison across all methods.
@@ 4.2 Main Results
Performance Comparison
Table 1 shows that our proposed ChronoRAG outperforms all baselines on the NarrativeQA dataset, with particularly strong gains on the "Time Question" subset. An analysis of the baselines on this dataset reveals distinct failure modes corresponding to their retrieval strategies. Summarization-based methods like RAPTOR-CT are the next-best performers but still lag our method; RAPTOR’s approach of clustering semantically similar chunks is insufficient for narratives, as it can group thematically related but chronologically distant events. Similarly, methods retrieving isolated text units, such as NaiveRAG and Propositionizer, struggle to provide sufficient context and fragment the narrative flow. GraphRAG records the lowest score, as its exhaustive entity–relation extraction adds thousands of trivial nodes, burying key plot elements under noise and severely diluting precision.
However, on the GutenQA dataset, the results are more nuanced. We explain that this is likely due to the nature of questions in GutenQA, which are often constructed by extracting text directly from the source and thus reward methods with the most direct access to the original passage details. For this reason, methods that heavily refine or abstract the text, such as Propositionizer and the graph-based approaches, perform poorly because they lose the specific passage-level information required. Even RAPTOR is penalized, as its summary-first approach limits direct access to the source text. In contrast, both NaiveRAG and ChronoRAG employ a consistent, fixed-length chunking strategy, which helps normalize the potentially unrefined structure of the GutenQA data. NaiveRAG’s slight edge in some metrics can be attributed to its direct retrieval from these unaltered chunks, whereas ChronoRAG’s abstraction steps, while beneficial for narrative synthesis, risk filtering out the fine-grained details that these specific questions demand.
Figure 4: A qualitative comparison of retrieved context and model answers for the query, "Where is George Darrow residing when he prepares to join Anna Leath in France?".
Ablation Study
We conduct ablation studies to investigate the effectiveness of different components and settings of ChronoRAG. As shown in the Table 2, without summarizing the original text and extracting entity relations shows a significant performance degradation, showing the importance of chunk summarization. The effects of summarization are twofold: it leaves only important information, making retrieval easier, and when assembling, it clarifies the flow. The results without passage assembling are obtained by individually searching for entity relations extracted from the summary, while the results without chunk summarization are obtained by searching for entity relations directly extracted from the 10 chunks. Despite not connecting nearby passages in both settings, a significant performance difference is observed in the TimeQuestion.
@@
Table 2: Ablation study of ChronoRAG’s core components on the NarrativeQA dataset, with performance measured by ROUGE-L.
4.3 Analysis
Trade-off between Linking Window and the Number of Retrieved Passage
We analyze the trade-off of using a larger linking window for passage assembly. While a wider window provides more local context, it also reduces the number of distinct passages that can be retrieved within a fixed token budget. Our experiment confirms this is detrimental; as shown in Table 3, extending the window to two neighbors ("Extended Link Window") lowers performance on both the whole dataset and the temporal questions subset. This result validates that our default approach of using a more concise, immediately adjacent context is more effective.
NarrativeQA Whole Data Time Question
ChronoRAG 0.308 0.268
Extended Link Window 0.300 0.258
Merged Key 0.302 0.257
Table 3: Analysis of design variations within the ChronoRAG framework on the NarrativeQA dataset measured by ROUGE-L.
Key-Value Separation in Information Retrieval
We investigate the effectiveness of key-value retrieval design of ChronoRAG, which separates the precise fact used for retrieval (the key) from the broader context provided to the model (the value). To validate this, we test an alternative "Merged Key" approach where the retrieved fact and its neighbors are combined into a single text unit before retrieval. As shown in Table 3, this modification results in a slight performance decrease, indicating that our key-value separation is an effective strategy for balancing retrieval precision and contextual coherence.
Effect of Neighborhood Assembling
# of Sentence mean(Similarity)
Retrieved Sentences 104,981 0.838
Assembled Sentences 209,092 0.785
Table 4: Embedding Similarity Between Query and Retrieved/Assembled Passages.
We validate the effectiveness of neighborhood assembling in enriching the retrieved context with information that is chronologically relevant but not necessarily the most semantically similar to the query. To demonstrate this, we compare the average embedding similarity between the query and the initially retrieved sentences versus the final assembled passages. As shown in Table 4, the average similarity for the assembled passages is discernibly lower than that of the sentences retrieved by similarity alone. This gap indicates that the neighboring passages, while chronologically adjacent, are semantically distinct from the initial query hit. In narrative texts where surrounding content carries strong causal or temporal relevance, this mechanism allows the model to incorporate pertinent information beyond the limits of pure similarity search, thereby improving the context for answer generation.
Case Study